OpenStudy (anonymous):

Given f(x) = 4x+1/3 solve for f^-1(3). A. 2 B. 4 C. 6 D. 8

OpenStudy (anonymous):

@phi

OpenStudy (phi):

is this \[ f(x) = \frac{4x+1}{3} \]?

OpenStudy (anonymous):

Yes! @phi

OpenStudy (phi):

this part \( f^{-1}(3) \) means the want the inverse function evaluated at 3. or, what x makes this true: \[ f(x) = \frac{4x+1}{3} \\3 = \frac{4x+1}{3} \] can you find the x that makes it true ?

OpenStudy (anonymous):

Makes it true? I don't understand...Where does the three come in to play? How do I solve? :O

OpenStudy (anonymous):

Wait! Would it be 2

OpenStudy (anonymous):

Because when you plug in 2 to the equation it comes to 3=3

OpenStudy (phi):

yes, the answer is 2

OpenStudy (anonymous):

A positive AP cal AP alg AP mR